# DeePoly: A High-Order Accuracy and Efficiency Deep-Polynomial Framework for Scientific Machine Learning

### v1

@@ 1 Introduction
Deep learning, empowered by the remarkable non-linear approximation capabilities of deep neural networks (DNNs) [5, 3], has emerged as a transformative paradigm in scientific computing. Among various deep learning approaches, Physics-Informed Neural Networks (PINNs) [19] have attracted substantial attention across engineering disciplines as an innovative methodology for solving Partial Differential Equations (PDEs). These methods offer promising alternatives to traditional numerical approaches such as the Finite Element Method (FEM) and Finite Difference Method (FDM) [16], particularly when addressing challenges that conventional methods struggle with. The widespread interest in PINNs stems from their distinctive advantages: they operate without requiring explicit mesh generation, eliminate the need for discretization schemes, and demonstrate remarkable versatility in handling both forward and inverse problems [19]. Despite these compelling features, PINN-based approaches encounter significant challenges in achieving the error convergence rates and computational efficiency characteristic of established numerical methods [22]. Quantitative studies have shown that even for relatively simple benchmark problems, PINNs typically require orders of magnitude more computational resources while achieving lower accuracy compared to traditional numerical solvers [10, 23]. These limitations, which arise from fundamental theoretical constraints in DNN approximation theory [15] and the inherent difficulties of non-convex optimization [26], have substantially restricted the broader adoption of PINNs in practical engineering applications.
@@
For accuracy enhancement, several strategies have emerged: (1) Advanced optimization techniques, including adaptive learning rate methods and second-order optimizers [11, 2], which can improve solution accuracy but often at significantly increased computational cost; (2) Modified network architectures [22, 9, 23] and high-frequency feature incorporation [21, 25], which can enhance fitting accuracy for solutions containing high-frequency structures, though typically only elevating challenging problems to accuracy levels comparable with low-frequency problems; and (3) Multi-stage neural network methods [24], which iteratively approximate residuals and have demonstrated machine-precision accuracy but may face limitations in efficiency and generalization to complex non-linear systems.
For efficiency improvement, approaches include: (1) Shallow random networks such as Extreme Learning Machines (ELM) [7, 6] and Random Feature Models (RFM) [17, 18], which can simultaneously improve accuracy and efficiency for smooth problems but often exhibit insufficient approximation capabilities for complex problems with large gradients or discontinuities; (2) Domain decomposition techniques [12, 8], which parallelize computation across subdomains but introduce challenges in maintaining solution continuity across boundaries; and (3) Alternative network formulations like Kolmogorov-Arnold Networks (KAN) [13, 1], which offer promising theoretical properties but remain in early stages and didn’t show accuracy improvement than PINNs in PDE applications.
@@
From an optimization perspective, this paradigm fundamentally transforms the solution process from relying solely on challenging non-convex parameter optimization to a more mathematically tractable approach: finding an acceptable local optimum through parameter optimization, followed by refinement through linear space optimization. Theoretical analysis and numerical experiments demonstrate that this approach not only significantly enhances both accuracy and efficiency compared to pure DNN methods but also inherits the well-established convergence properties of polynomial approximation while maintaining the key advantages of PINNs, such as being mesh-free and scheme-free.
@@ 2.1 From Nonlinear Optimization to Nonlinear Features Extraction and Linear Space Optimization
Consider a typical non-convex optimization problem as illustrated in Figure 1. When parameter optimization begins from point A, conventional approaches can readily find nearby local minima such as points B or C. However, locating smaller local minima or the global minimum remains exceptionally challenging due to two fundamental constraints:
  1. 1.
    The smaller local minima may be separated by substantial distances in parameter space or potentially non-existent.
  2. 2.
    Probabilistically, in complex landscapes with numerous local minima, smaller local minima tend to be sparser.
@@
Figure 1: Illustration of the New Method
@@
Step 1 (Spotter): Through parameter optimization, we find an acceptable solution with moderate error (for example, $Loss=10^{-3}$), denoted as $NN(\theta_{B};x)$. At this point, we cease parameter optimization, avoiding the diminishing returns associated with pursuing lower local minima through non-convex optimization.
Step 2 (Sniper): We extract the network features from the trained network at point B as rough global features:
$NN(\theta_{B};x)=\sum_{n=1}^{N}\beta_{n}\xi_{n}(x),\quad x\in X$
@@
$\mathbf{W}=\text{span}\{\xi_{1}(x),\dots,\xi_{N}(x)\}.$
Then, we supplement these features with polynomial basis functions to address the low-order features potentially missed by the neural network (as elaborated in the Theoretical section 4):
$\mathbb{P}=\text{span}\{P_{0}(x),P_{1}(x),\dots,P_{M}(x)\}$
where $P_{i}(x)$ are polynomial basis functions. We construct the combined linear space:
$\mathbf{V}=\mathbf{W}\oplus\mathbb{P}$
We then perform linear optimization within the space $\mathbf{V}$ to find the best approximation, aiming to reach point E:
$\min_{\alpha_{n}\in\mathbb{R}}\left\|\sum_{n=1}^{K}\alpha_{n}v_{n}(x)-t(x)\right\|\quad\text{where}\quad v_{n}(x)\in\mathbf{V},K=\dim(\mathbf{V})$
@@
Since continuous low-order polynomial features possess controllable spatial convergence properties (see Theoretical section), if point E does not meet the required accuracy, we can employ domain partitioning. The Sniper optimization is then performed on the combination of global features and local low-order features within smaller subdomains, potentially reaching machine precision.
@@ 2.2 Solving Time-Dependent Nonlinear PDEs
With the assistance of automatic differentiation, the solution process for PDEs follows a similar approach to function approximation, with special consideration given to the linearization of non-linear terms. While the optimization process described above is effective for directly solving non-linear functions—converting from fine-tuning network coefficients to fine-tuning fewer linear coefficients—we recommend linearizing non-linear terms to guarantee accuracy, given the challenges of non-linear optimization. Time discretization and pseudo-time stepping represent two effective approaches for this linearization.
@@ 3.3.2 Linear Convection Equation
It should be noted that while the current accuracy for discontinuous problems does not reach the machine precision levels achieved for smooth problems, this limitation stems from the conditioning properties of the large linear system that must be solved in the presence of discontinuities. Future work will explore the integration of Physics-Informed Neural Networks with Weighted Essentials (PINNs-WE) [14] to effectively modulate the influence of discontinuous regions on the matrix conditioning, thereby further enhancing accuracy in smooth regions.

### v2

# DeePoly: A High-Order Accuracy Scientific Machine Learning Framework for Function Approximation and Solving PDEs

@@ 1 Introduction
Deep learning, empowered by the remarkable non-linear approximation capabilities of deep neural networks (DNNs) [9, 6], has emerged as a transformative paradigm in scientific computing. Among various deep learning approaches, Physics-Informed Neural Networks (PINNs) [22] have attracted substantial attention across engineering disciplines as an innovative methodology for solving Partial Differential Equations (PDEs). These methods offer promising alternatives to traditional numerical approaches such as the Finite Element Method (FEM) and Finite Difference Method (FDM) [21], particularly when addressing challenges that conventional methods struggle with. The widespread interest in PINNs stems from their distinctive advantages: they operate without requiring explicit mesh generation, eliminate the need for discretization schemes, and demonstrate remarkable versatility in handling both forward and inverse problems [22]. Despite these compelling features, PINN-based approaches encounter significant challenges in achieving the error convergence rates and computational efficiency characteristic of established numerical methods [14, 28]. Quantitative studies have shown that even for relatively simple benchmark problems, PINNs typically require orders of magnitude more computational resources while achieving lower accuracy compared to traditional numerical solvers [14, 28]. These limitations, which arise from fundamental theoretical constraints in DNN approximation theory [20] and the inherent difficulties of non-convex optimization [31], have substantially restricted the broader adoption of PINNs in practical engineering applications.
@@
For accuracy enhancement, several strategies have emerged: (1) Advanced optimization techniques, including adaptive learning rate methods and second-order optimizers [16, 2, 26], which can improve solution accuracy but often at significantly increased computational cost; (2) Modified network architectures [27, 13, 28] and high-frequency feature incorporation [25, 30], which can enhance fitting accuracy for solutions containing high-frequency structures, though typically only elevating challenging problems to accuracy levels comparable with low-frequency problems; and (3) Multi-stage neural network methods [29], which iteratively approximate residuals and have demonstrated machine-precision accuracy but may face limitations in efficiency and generalization to complex non-linear systems.
For efficiency improvement, approaches include: (1) Shallow random networks such as Extreme Learning Machines (ELM) [11, 10, 7] and Random Feature Models (RFM) [4], which can simultaneously improve accuracy and efficiency for smooth problems but often exhibit insufficient approximation capabilities for complex problems especially with large gradients or discontinuities; (2) Domain decomposition techniques [17, 12], which parallelize computation across subdomains but introduce challenges in maintaining solution continuity across boundaries; and (3) Alternative network formulations like Kolmogorov-Arnold Networks (KAN) [18], which offer promising theoretical properties but remain in early stages and didn’t show accuracy improvement than PINNs in PDE applications [23].
@@
From an optimization perspective for DNNs, research inspired by the p-spin glass model in physics [3, 1] reveals that their loss landscapes are largely free of "poor" local minima [5, 15]. In other words, the performance gap between most local minima and the global optimum is small, which allows algorithms to efficiently find a good solution near the initial parameters. However, this negligible difference from the perspective of network training may prove fatal in scientific computing—for instance, the distinction between $10^{-3}$ and $10^{-5}$ errors. Such fundamental divergence in precision tolerance across domains compels us to seek better local optima. The true challenge lies in the fact that while these high-quality local minima are virtually identical in performance, they may lie extremely far apart in parameter space. Consequently, any attempt to explore between these distant solutions to extract marginal performance gains inevitably leads to a dramatic surge in training costs. Based on the above rationale, this paradigm fundamentally transforms the solution process: rather than relying solely on challenging non-convex parameter optimization, it adopts a mathematically tractable two-stage approach. First, the Scoper step efficiently identifies an acceptable solution near the initial parameters-achieving potentially substantial loss reduction (without necessarily converging to local minima) through computationally lightweight optimization. Second, the Sniper step employs high-order polynomial bases to refine accuracy via linear subspace optimization. Theoretical analysis and numerical experiments demonstrate that this framework not only significantly enhances both accuracy and efficiency compared to pure DNN methods but also inherits the rigorous convergence guarantees of polynomial approximation while preserving key advantages of Physics-Informed Neural Networks (PINNs), including broad applicability, stability, mesh-free operation, and scheme independence.
@@ 2.1 From Nonlinear Optimization to Nonlinear Features Extraction and Linear Space Optimization
Consider a typical non-convex optimization problem as illustrated in Figure 1.(a-b). When parameter optimization begins from point A, conventional approaches can readily find nearby local minima such as points B. However, locating smaller local minima or the global minimum remains exceptionally challenging and expensive.
@@
Figure 1: Graphic illustration of DeePoly
(a) The local minimas are close to the global minimum however they are separated in parameter space. (b) It is effective to find a relatively good local minima however it expensive to traverse from one local minima to another. (c) Scoper works as a features extractor to obtain similar features of the target, without the requirement of high-precision approximation (d) Sniper works locally as a corrector to gain high-order accuracy. (e) Theoretically, DeePoly can achieve high-order (same to the order of the polynomial basis) accuracy by the guarantee of Taylor expansion of smooth function, and the requirement of partitions is lower than traditional methods.
@@
Step 1 (Spotter): Through parameter optimization, we find an acceptable solution with moderate error as illustrated as Fig.1.(a-c) (for example, $Loss=10^{-3}$), denoted as $f_{\rm Scoper}(\theta_{B};x)$. At this point, we cease parameter optimization, avoiding the diminishing returns associated with pursuing lower local minima through non-convex optimization.
Step 2 (Sniper): We extract the network features from the trained network at point B as approximate global features:
$f_{\rm Scoper}(\theta_{B};x)=\sum_{n=1}^{N}\beta_{n}\xi_{n}(x),\quad x\in\Omega$
@@
$\boldsymbol{\xi}=\text{span}\{\xi_{1}(x),\dots,\xi_{N}(x)\}.$
Then we supplement these features with polynomial basis functions to gain high-order accuracy with partition (as elaborated in the Theoretical section 4):
$\boldsymbol{P}=\text{span}\{P_{0}(\bar{x}),P_{1}(\bar{x}),\dots,P_{M}(\bar{x})\}$
where $P_{i}(\bar{x})$ are polynomial basis functions and $\bar{x}\in\Omega_{j}$ is the partition of the domain $\Omega$,and
$\bar{x}=\frac{x-\inf(\Omega_{j})}{\sup(\Omega_{j})-\inf(\Omega_{j})}$
Then we construct the combined linear space:
$\mathbf{V}=\boldsymbol{\xi}\oplus\boldsymbol{P}$
and perform linear optimization within the space $\mathbf{V}$ to find the best approximation, aiming to reach point E:
$\min_{\gamma_{n}^{i}\in\mathbb{R}}\left\|\sum_{n=1}^{K}\gamma_{n}^{i}v_{n}(x)-t(x)\right\|\quad\text{where}\quad v_{n}(x)\in\mathbf{V},K=\dim(\mathbf{V})\quad x\in\Omega_{i}$
@@
Since continuous low-order polynomial features possess controllable spatial convergence properties (see Theoretical section), if point E does not meet the required accuracy, we can employ domain partitioning. The Sniper optimization is then performed on the combination of global features and local M-order features within smaller subdomains, potentially reaching machine precision theoretically.
@@ 2.2 Solving Time-Dependent Nonlinear PDEs
With the assistance of automatic differentiation, the solution process for PDEs follows a similar approach to function approximation, with special consideration given to the linearization of non-linear terms. While the optimization process described above is effective for directly solving non-linear functions-converting from fine-tuning network coefficients to fine-tuning fewer linear coefficients-we recommend linearizing non-linear terms to guarantee accuracy, given the challenges of non-linear optimization. Time discretization and pseudo-time stepping represent two effective approaches for this linearization.
@@ 3.3.2 Linear Convection Equation
It should be noted that while the current accuracy for discontinuous problems does not reach the machine precision levels achieved for smooth problems, this limitation stems from the conditioning properties of the large linear system that must be solved in the presence of discontinuities. Future work will explore the integration of Physics-Informed Neural Networks with Weighted Essentials (PINNs-WE) [19] to effectively modulate the influence of discontinuous regions on the matrix conditioning, thereby further enhancing accuracy in smooth regions of discontinuous problems [19].